Assume that 10 ft-lb of work is required to stretch a spring 1 ft beyond its natural length. What is the spring constant?


1
Expert's answer
2022-05-03T13:58:04-0400

"W=\int^b_a kxdx=k\int^b_axdx=k(x^2/2)|^b_a =k(b^2/2-a^2/2)"

W=10

a=0

b=1

"10=k(1\/2-0)=k\/2"

k=20